Draw FCC structure By considering the packing ofplease answer b,c,d
Answer #1

FCC Coordination number is the number of nearest neighbor to a particular atom in the crystal coordination number 12 total 4

Atomic packing factor for FCC

Let a be the A the side length of the unit cell of FCC lattice and R the diameter of the atoms.

The FCC unit cell is formed by 4 atoms:
- 8 times one eighth of an atom at the corners of the cube
- 4 times a half of an atom at the center of the of the faces.

At the faces the atoms at the corners and the center atom touch, so that the perfectly fill the face. Hence the length of the face diagonal is
D = R + 2R + R = 4R
From Pythagorean theorem you get
A² + A² = D²
=>
A = √8 · R = √2 · 2·R

The volume of the cube cell is
Vc = A³ = √2 · 16·R
The volume of the atoms in the cell is
Va = 4 · (4·π·R³ /3) = 16·π·R³ /3

The packing density is
p = Va / Vc
= (16·π·R³ /3) / (√2 · 16·R)
= π / (3·√2)
= 0.74